# Using Enriched Category Theory to Construct the Nearest Neighbour Classification Algorithm

### v1

@@ Abstract
Exploring whether Enriched Category Theory could provide the foundation of an alternative approach to Machine Learning. This paper is the first to construct and motivate a Machine Learning algorithm solely with Enriched Category Theory. In order to supplement evidence that Category Theory can be used to motivate robust and explainable algorithms, it is shown that a series of reasonable assumptions about a dataset lead to the construction of the Nearest Neighbours Algorithm. In particular, as an extension of the original dataset using profunctors in the category of Lawvere metric spaces. This leads to a definition of an Enriched Nearest Neighbours Algorithm, which consequently also produces an enriched form of the Voronoi diagram. This paper is intended to be accessible without any knowledge of Category Theory.
Keywords: Category Theory, Enriched Category Theory, Lawvere Metric Space, Classification, Nearest Neighbours, Foundations
@@
As Machine Learning (ML) becomes more popular, the use of black box approaches is beginning to hinder the progression of the field. During engineering and development, the better ones understanding of a model the easier it is to improve its performance, diagnose faults, and provide guarantees for its behaviour. Unfortunately, necessary to the development of many algorithms, there are design decisions which are motivated by intuition or trial and error. Potentially, part of the difficulty in understanding these algorithms comes from a lack of clarity in how they are interacting with the data they are provided. How does the encoding of input data effect the information that an algorithm actually understands. To approach this question, this paper seeks to investigate the development of a first principles approach to the design of ML algorithms using Enriched Category Theory. To provide evidence that this approach has potential, it is demonstrated that basic assumptions about a dataset can lead to the natural construction of a pre-existing algorithm which is popular for its predictable and robust behaviour: the Nearest Neighbours Algorithm (NNA).
The argument for the use of Enriched Category Theory in such a theory proceeds as follows. The process of learning requires the ability to make comparisons. This may be comparisons between: entries of a training dataset in order to identify patterns; training examples and new cases for the sake of inference; between different models of the same dataset, for selection of the best one. Enriched Category Theory provides a very general framework for defining and studying comparisons between objects. It demonstrates that the entirety of the information associated with an object can be encoded in its comparisons to other objects. Using Enriched Category Theory, the structure of data can be encoded explicitly in their mutual comparisons, rather than implicitly, as is common with many ML algorithms. The benefit of this approach would be that the design and mechanism of ML algorithms would becomes more transparent. The assumptions about datasets can be made more explicit. And the process of learning can be interpretted in its natural form as reasoning about the comparison of observations.
@@ 2 Background
To the knowledge of the authors, the construction of the Nearest Nearest Neighbours Algorithm demonstrated in this paper is one of the first examples of a machine learning algorithm motivated and constructed solely with Enriched Category Theory. There is one other example of an entirely categorical construction of an ML algorithm, where Shiebler (2022) shows that the single linkage clustering algorithm can be found as a Kan-extension of a dataset of points. However, it is suggested that the steps shown for the derivation of the NNA draw a tighter parallel between the intuition of how the dataset is represented, and the derived algorithm.
There are also examples of algorithms whose structures have been encoded in the language of category theory, such as Graph Neural Networks (Dudzik and Veličković, 2022). But they represent the structure of how the algorithm computes information, and not necessarily the selection of the optimal model or representation of the input dataset. In contrast, the NNA construction draws a direct line from the representation of the data to the selection of the optimal classification.
Understanding the Enriched Category Theory construction of the Nearest Neighbours algorithm requires an understanding of Lawvere metric space as Cost Enriched Categories, as well as a working knowledge of the Nearest Neighbours Algorithm. It is beyond the scope of this paper to provide a complete introduction to Enriched Category Theory  [^1] A basic introduction can be found in Fong and Spivak (2018) while a more technical overview occurs in Kelly (2005)., but thankfully many of its complexities can be avoided by focusing on the specific case of Lawvere metric spaces. The following section provides the necessary components, as well as a brief overview of the Nearest Neighbours Algorithm.
@@ 2.1 Nearest Neighbours Algorithm
The Nearest Neighbours Algorithm (Fix and Hodges, 1989) extends the classification of a dataset of points in a metric space to the entire metric space. Consider a dataset of n pairs $(x_{0},y_{0}),...,(x_{n},y_{n})$. The targets of the dataset, $y_{i}$, are elements of a set of class labels $Y$. The features of the dataset, $x_{i}$, represent points in a metric space $X$. This allows the distance between any two points to be measured, following the traditional metric space axioms.
@@
To a point of the metric space not in the dataset, the Nearest Neighbors Algorithm assigns a class if the closest point in the dataset has that class. An example of the classification regions produced can be seen in Fig 1 which shows the NNA classification of a two class dataset of points sampled from two Gaussian distributions.
@@
$NNA(y,x)=\exists i\in N[\quad Ti=y\quad and\quad d(Fi,x)=\inf_{i^{\prime}\in N}d(Fi^{\prime},x)\quad]$
This relation can be presented in an alternate form that will be useful later, but it requires that the indexes are partitioned based on their classes. We define the partition as follows. $NT(y)=\{i\in N\ |\ Ti=y\}$. This allows the relation to be presented as:
$NNA(y,x)\quad\Leftrightarrow\quad\inf_{i\in N}d(Fi,x)\quad=\inf_{i\in NT(y)}d(Fi,x)$
@@ 2.2 Lawvere Metric Spaces
As mentioned in the introduction, Enriched Category Theory provides a method of encoding structure through a rigorous language for talking about comparisons. In some sense, an Enriched Category is a collection of objects which can be compared. Given a category $C$, two objects $x\in C$ and $y\in C$ can be compared with the notation $C(x,y)$. This is referred to as the hom-object of $x$ and $y$. This hom-object exists in its own category called the base of enrichment. To make the comparisons meaningful, ECT requires that the base of enrichment have some way of combining hom-objects, called a monoidal product, and some juxtaposition of these two hom-objects to a third. An example of how this structure works can be seen in order relations. Consider a category called Fruits, which is a collection of fruits ordered by price. The hom-object $Fruits(Apple,Orange)$ would test to see if Apples were cheaper than Oranges. In this instance this comparison could also be written as $Apples\leq Oranges$. The outcome of this comparison is either true or false so the base of enrichment would be a category containing an object representing true and an object representing false. This base of enrichment can be called Bool for Boolean.
@@
A sensible logical deduction to make with such a category would be to say that if I know fruit $A$ is cheaper than fruit $B$, and fruit $B$ is cheaper than fruit $C$, then $A$ must be cheaper than $C$. Notionally, this can be written as:
$(A\leq B)\ and\ (B\leq C)\implies(A\leq C)$
@@
This process of logical inference gives the general motivating structure of an enriched category. In this instance, each comparison of the ordered set returns a value in Bool. The monoidal product of Bool is the logical ”and”, allowing its objects to be combined. Bool also has arrows of implication from False to False, False to True, and True to True. But not from True to False, as True cannot logically imply False. By using Bool as the base of enrichment, the general structure of the enriched category becomes the structure of a pre-order relation.
@@
A Lawvere metric space is an enriched category whose base of enrichment is chosen so that the categories operate like metric spaces, allowing the enriched category to measure the distances between its objects. The base of enrichment for a Lawvere metric spaces is called the Cost category. Because it represents measurements of distance, its objects are the non-negative real numbers extended with infinity [^3] The objects of Cost being $\{x\in\mathbb{R}\ |\ x\geq 0\}\cup\{\infty\}$. The monoidal product is addition, with addition by infinity defined as $x+\infty=\infty$.Given a Cost enriched category $X$, and two objects $x$ and $y$ of $X$, the hom object $X(x,y)$ can be interpretted as the distance between $x$ and $y$. The monoidal product of Cost is addition and the arrows of the Cost category point from large numbers to smaller numbers. As in, there is an arrow from $a\in Cost$ to $b\in Cost$ if and only if $a\geq b$. This can also be interpretted as $Cost(a,b)=a\geq b$. Looking at the previous example, we can replace the $and$ operation of Bool with addition, and the implication with $\geq$ to recover the following expression for Cost categories.
$X(x,y)+X(y,z)\geq X(x,z)$
@@
This requirement of Cost categories is the triangle inequality, stating that taking a detour to a third object cannot be quicker than travelling directly between two objects. By choosing the Cost category as the base of enrichment, ECT naturally recovers some, but not all of the the metric space axioms (As detailed in section 2.1). This makes Lawvere metric spaces pseudo-metric spaces. In Lawvere metric spaces, one retains the triangle inequality, and the requirement that the distance from an object to itself is zero ($d(a,a)=0$), but the metric spaces are not required to be symmetric ($d(a,b)=d(b,a)$) and two different objects can be zero distance apart. This can be a controversial choice, but there are several arguments for this being a desirable outcome. For example, in many cases an intuitive notion of distance is not symmetric, e.g. its easier to go down stairs than up them. One might also say that distance is a measure of similarity not identity, and the idea of two different objects being zero distance apart is sensible when considering systems at a certain level of coarseness. In either case, if one wishes to operate with traditional metric spaces, they are all also Lawvere metric spaces, and the necessary axioms can be asserted as convenient.
@@
By sensibly considering how we wish to compare objects in our enriched categories, choosing objects, arrows, and a monoidal product in the base of enrichment, we have recovered the structure of a metric space. Though the Lawvere metric space is one of the simpler examples of an enriched category, it starts to reveal the power of such a theory to construct complex structures for the representation of data.
@@ 2.3 Functors and Profunctors
An Enriched Category may be thought of as representing a particular datatype, with the structure of that datatype being represented by the hom-objects of the category. In order to interact with this information, there are many ways of comparing categories to each other. Between categories with the same base of enrichment, there are two constructions which are relevant for this work: Functors and Profunctors.
@@
$X(a,b)\geq Y(Fa,Fb)$
As well as Functors between categories being the ECT version of functions between sets, there is also an ECT version of relations between categories. A set relation $R$ between two sets $X$ and $Y$ is often described as a subset of the Cartesian products of $X$ and $Y$, i.e. $R\subseteq X\times Y$. However, this relation can also be thought of as a function which returns true if the relation is true, and false if the relation is false: $R:Y\times X\rightarrow\{False,True\}$. In ECT, this notion is extended to a functor from the product of two categories to the base of enrichment. Where the product of two categories $Y^{op}\otimes X$ contains objects which are pairs of objects in $X$ and $Y$ similar to how the Cartesian product of sets contains pairs of elements of sets. $R:Y^{op}\otimes X\rightarrow Cost$. Such a construction is called a profunctor. For notation, a profunctor $R:Y^{op}\otimes X\rightarrow Cost$, can be written as $R:X\nrightarrow Y$.
@@
$(S\circ R)(z,x):=\inf_{y\in Y}\ (R(y,x)+S(z,y))$
The similarity between relation composition and profunctor composition is more than just cosmetic. It also emulates how Cost enriched categories treat logical propositions. In the Boolean logic setting, the ”and” operation outputs true only when both of its inputs are true, and false otherwise. In Cost enriched categories, a distance of zero can be interpreted as true, and a distance greater than zero is false. With this interpretation, the sum of two values $a$ and $b$, where both are non-negative, can only be zero if both $a$ and $b$ are zero. From the perspective of Cost category logic, $a+b$ is the logical ”and” operation. Furthermore, within this version of logic the infimum operation is the Cost version of the existential quantifier. When X is finite, The statement $\inf_{x\in X}Fx=0$ means there exists a value $x$ such that $Fx$ is zero. In the infinite case, it suggests that there exists a value $Fx$ which is arbitrarily close to zero. Applying this logic to the definition of profunctors, it can be seen that profunctors produce truth values from pairs of objects, if the output of zero is interpreted as true, and the output of non-zero is interpretted as false. Such an interpretation can be represented by the functor $(0=x):Cost\rightarrow Bool$.
With knowledge of Functors, Profunctors and their composition there is a final piece of information necessary for the construction of the Nearest Neighbours Algorithm. Continuing with the intuition from functions and relations of sets, it can be observed that functions are a special kind of relation, known as a functional relation. A function $F:N\rightarrow X$ is said to produce an element $Fi$ when given an element $i\in N$, but this behaviour can be represented directly as a relation $F_{*}:N\nrightarrow X$ which evaluates to a truth value under the condition $F_{*}(x,i)\Leftrightarrow(x=Fi)$. In fact, there is also a second relation of the opposite direction $F^{*}:X\nrightarrow N$ which represents the logical evaluation of the function $F^{*}(i,x)\Leftrightarrow(Fi=x)$.
The interaction between functions and relations has a mirror in the interaction between functors and profunctors. A functor $F:N\rightarrow X$ canonically generates two profunctors. One of the same direction $F_{*}:N\nrightarrow X$ and one of the opposite direction $F^{*}:X\nrightarrow N$. They are defined with the aid of hom-objects, where $F_{*}(x,i)=X(x,Fi)$ and $F^{*}(i,x)=X(Fi,x)$. In the case of Lawvere metric spaces, the profunctors of $F$ evaluated on objects $x$ and $i$ can be read as: ”The distance between $x$ and the image of $i$ under $F$”. With this final component, it is now possible to construct the Nearest Neighbours Algorithm.

### v2

@@ Abstract
This paper is the first to construct and motivate a Machine Learning algorithm solely with Enriched Category Theory, supplementing evidence that Category Theory can provide valuable insights into the construction and explainability of Machine Learning algorithms. It is shown that a series of reasonable assumptions about a dataset lead to the construction of the Nearest Neighbours Algorithm. This construction is produced as an extension of the original dataset using profunctors in the category of Lawvere metric spaces, leading to a definition of an Enriched Nearest Neighbours Algorithm, which, consequently, also produces an enriched form of the Voronoi diagram. Further investigation of the generalisations this construction induces demonstrates how the $k$ Nearest Neighbours Algorithm may also be produced. Moreover, how the new construction allows metrics on the classification labels to inform the outputs of the Enriched Nearest Neighbour Algorithm: Enabling soft classification boundaries and dependent classifications. This paper is intended to be accessible without any knowledge of Category Theory.
Machine Learning, polynomial, optimisation theory
@@ 1 Introduction
Recent work has demonstrated that category theory may be useful in the analysis of machine learning algorithms(Shiebler et al., 2021). However, there are almost no examples of a category theoretic construction being used for both the model description and selection. Such a construction is necessary to allow future researchers to apply the tools of category theory in the analysis of machine learning algorithms.
This work has produced a construction of the Nearest Neighbours Algorithm (NNA),which describes both its modelling behaviour and selection entirely within the language of Enriched Category Theory (ECT). From this initial result, a series of observations about the structure of the algorithm can be made, which give an indication of how category theory may be used in the analysis of machine learning algorithms.
Firstly, the construction of the NNA uses a category called Cost as its base of enrichment, which produces enriched categories that operate like metric spaces. This allows standard intuitions about the NNA to inform the construction. However, no part of the construction depends on Cost being the base of enrichment. This makes the first outcome the production of an infinite family of algorithms for any choice of enrichment base, called V-NNA.
Secondly it is shown that the k-Nearest Neighbours Algorithm ($k$-NNA) can be constructed using the categorical description of the NNA. It demonstrates that any $k$-NNA may be presented as a NNA when the data is projected into a metric space of tuples.
Thirdly, the construction of the NNA allows not only the data points to inhabit a metric space, but also their class labels. This causes the construction of the NNA to generalise the algorithms behaviour such that it may operate with soft or dependent classifications.
Finally, the category of cost enriched categories which the NNA lives inside of is equipped with an internal language which may express concepts using distance, addition, truncated subtraction, infimum, and supremum. Because both model representation and selection occurs inside this category, the NNA is only capable of understanding information expressible in these terms. The relationship between the internal language of the category and the behaviour of the NNA has interesting implications for algorithm explainability.
@@ 2 Background
The construction of the Nearest Neighbours Algorithm demonstrated is the first examples of a machine learning algorithm constructed with Enriched Category Theory (ECT). There is one other example of an entirely categorical construction of an ML algorithm, where previous work (Shiebler, 2022) shows that the single linkage clustering algorithm can be found as a Kan-extension of a dataset of points.
There are examples of algorithms whose structures have been encoded in the language of category theory, such as Graph Neural Networks (Dudzik & Veličković, 2022). But they do not represent how the algorithm selects the best model.
It is beyond the scope of this paper to provide a complete introduction to Enriched Category Theory [^1] A basic introduction can be found in ”Seven Sketches in Compositionality” (Fong & Spivak, 2018) while a more technical overview occurs in ”Basic Concepts of Enriched Category Theory” (Kelly, 2005)., but thankfully many of its complexities can be avoided by focusing on the specific case of Lawvere metric spaces.
@@ 2.1 Nearest Neighbours Algorithm
The Nearest Neighbours Algorithm (Fix & Hodges, 1989) extends the classification of a dataset of points in a metric space to the entire metric space. Consider a dataset of $n$ pairs $(x_{1},y_{1}),...,(x_{n},y_{n})$. The targets of the dataset, $y_{i}$, are elements of a set of class labels $Y$. The features of the dataset, $x_{i}$, represent points in a metric space $X$. This allows the distance between any two points to be measured, following the traditional metric space axioms.
@@
To a point of the metric space not in the dataset, the Nearest Neighbours Algorithm assigns a class if a closest point in the dataset has that class. An example of the classification regions produced can be seen in Fig 1 which shows the NNA classification of a two class dataset of points sampled from two Gaussian distributions.
@@
$\begin{split}NNA(y,x)=\exists i\in N\ [\quad&Ti=y\quad and\\ &\quad d(Fi,x)=\inf_{i^{\prime}\in N}d(Fi^{\prime},x)\quad]\end{split}$ (1)
Define the partition $NT(y)=\{i\in N\ |\ Ti=y\}$. This allows the relation to be presented as:
$NNA(y,x)\quad\Leftrightarrow\quad\inf_{i\in N}d(Fi,x)\quad=\inf_{i\in NT(y)}d(Fi,x)$ (2)
@@ 2.2 Lawvere Metric Spaces
Given an enriched category $C$, two objects $x\in C$ and $y\in C$ can be compared with the notation $C(x,y)$. This is referred to as the hom-object of $x$ and $y$. This hom-object exists in its own category called the base of enrichment. To make the comparisons meaningful, ECT requires that the base of enrichment have some way of combining hom-objects, called a monoidal product, and some juxtaposition of these two hom-objects to a third. An example of how this structure works can be seen in order relations. Consider a category called Fruits, which is a collection of fruits ordered by price. The hom-object $Fruits(Apple,Orange)$ would test to see if Apples were cheaper than Oranges. In this instance this comparison could also be written as $Apples\leq Oranges$. The outcome of this comparison is either true or false so the base of enrichment would be a category containing an object representing true and an object representing false. This base of enrichment can be called Bool for Boolean.
@@
A sensible logical deduction to make with such a category would be to say that if we know fruit $A$ is cheaper than fruit $B$, and fruit $B$ is cheaper than fruit $C$, then $A$ must be cheaper than $C$. Notionally, this can be written as:
$(A\leq B)\ and\ (B\leq C)\implies(A\leq C)$ (3)
@@
This process of logical inference gives the general motivating structure of an enriched category. In this instance, each comparison of the ordered set returns a value in Bool. The monoidal product of Bool is the logical ”and”, allowing its objects to be combined. The hom-object of Bool is logical implication.
$Bool(x,y)=(x\Rightarrow y)$ (4)
Because Bool can be described with a hom-object which takes values in itself, it is described as self enriched. When Bool is used as the base of enrichment, the general structure of the enriched category becomes the structure of a pre-order relation.
@@
A Lawvere metric space is an enriched category whose base of enrichment is chosen so that the categories operate like metric spaces, allowing the enriched category to measure the distances between its objects. The base of enrichment for Lawvere metric spaces is called the Cost category. Because it represents measurements of distance, its objects are the non-negative real numbers extended with infinity [^3] The objects of Cost being $\{x\in\mathbb{R}\ |\ x\geq 0\}\cup\{\infty\}$. The monoidal product is addition, with addition by infinity defined as $x+\infty=\infty$. Given a Cost enriched category $X$, and two objects $x$ and $y$ of $X$, the hom object $X(x,y)$ can be interpreted as the distance between $x$ and $y$. The monoidal product of Cost is addition. Cost can be interpreted as both a Bool enriched category, and a self enriched category. As a Bool enriched category, it is a preorder, with the ordering of its objects given by the standard order relation on the real numbers.
$Cost(x,y)=(x\geq y)$ (5)
Looking at the previous example, we can replace the $and$ operation of Bool with addition, and the implication with $\geq$ to recover the following expression for Cost categories.
$X(x,y)+X(y,z)\geq X(x,z)$ (6)
@@
This requirement of Cost categories is the triangle inequality, stating that taking a detour to a third object cannot be quicker than traveling directly between two objects. By choosing the Cost category as the base of enrichment, ECT naturally recovers some, but not all of the the metric space axioms (As detailed in section 2.1). This makes Lawvere metric spaces pseudo-metric spaces. In Lawvere metric spaces, one retains the triangle inequality, and the requirement that the distance from an object to itself is zero ($d(a,a)=0$), but the metric spaces are not required to be symmetric ($d(a,b)=d(b,a)$) and two different objects can be zero distance apart. This can be a controversial choice, but there are several arguments for this being a desirable outcome. For example, in many cases an intuitive notion of distance is not symmetric, e.g. its easier to go down stairs than up them. One might also say that distance is a measure of similarity not identity, and the idea of two different objects being zero distance apart is sensible when considering systems at a certain level of coarseness. In either case, if one wishes to operate with traditional metric spaces, they are all also Lawvere metric spaces, and the necessary axioms can be asserted as convenient.
@@
By sensibly considering how we wish to compare objects in our enriched categories, choosing objects and a monoidal product in the base of enrichment, we have recovered the structure of a metric space. Though the Lawvere metric space is one of the simpler examples of an enriched category, it starts to reveal the power of such a theory to construct complex structures for the representation of data.
@@ 2.3 Functors and Profunctors
$X(a,b)\geq Y(Fa,Fb)$ (7)
A set relation $R$ between two sets $X$ and $Y$ is often described as a subset of the Cartesian products of $X$ and $Y$, i.e. $R\subseteq X\times Y$. However, this relation can also be thought of as a function which returns true if the relation is true, and false if the relation is false: $R:Y\times X\rightarrow\{False,True\}$. In ECT , this notion is extended to a functor from the tensor product of two categories to the base of enrichment.
$R:Y^{op}\otimes X\rightarrow Cost$ (8)
Such a construction is called a profunctor. A profunctor $R:Y^{op}\otimes X\rightarrow Cost$, can be written as $R:X\nrightarrow Y$. The tensor product of two categories $Y^{op}\otimes X$ contains objects which are pairs of objects in $X$ and $Y$ similar to how the Cartesian product of sets contains pairs of elements of sets. The notation $Y^{op}$ is used to refer to a category with the objects of $Y$, but whose hom objects are reversed.
$Y(a,b)=Y^{op}(b,a)$ (9)
@@
$(S\circ R)(z,x):=\inf_{y\in Y}\ (R(y,x)+S(z,y))$ (11)
In Boolean logic, the ”and” operation outputs true only when both of its inputs are true, and false otherwise. In Cost enriched categories, a distance of zero can be interpreted as true, and a distance greater than zero is false. The sum of two values $a$ and $b$, where both are non-negative, can only be zero if both $a$ and $b$ are zero. In Cost category logic, $a+b$ is the logical ”and” operation. Furthermore, the infimum operation is the Cost version of the existential quantifier. When $X$ is finite, The statement $\inf_{x\in X}Fx=0$ means there exists a value $x$ such that $Fx$ is zero. In the infinite case, it suggests that there exists a value $Fx$ which is arbitrarily close to zero. Applying this logic to the definition of profunctors, it can be seen that profunctors produce truth values from pairs of objects. Such an interpretation can be represented by the functor $(0=x):Cost\rightarrow Bool$. The following table highlights the comparable notions between standard Set theory and Lawvere metric spaces, though this comparison extends to all ECT.
Logical Concept Set Theory Lawvere metric spaces
Truth Values Bool Cost
Conjunction $\wedge$ +
Mappings Functions Functors
Binary Predicate Relations Profunctors
Existential Quantifier $\exists$ $\inf$
Universal Quantifier $\forall$ $\sup$
It can be observed that functions are a special kind of relation, known as a functional relation. A function $F:N\rightarrow X$ is said to produce an element $Fi$ when given an element $i\in N$, but this behaviour can be represented directly as a relation $F_{*}:N\nrightarrow X$ which evaluates to a truth value under the condition $F_{*}(x,i)\Leftrightarrow(x=Fi)$. In fact, there is also a second relation of the opposite direction $F^{*}:X\nrightarrow N$ which represents the logical evaluation of the function $F^{*}(i,x)\Leftrightarrow(Fi=x)$.
The interaction between functions and relations has a mirror in the interaction between functors and profunctors. A functor $F:N\rightarrow X$ canonically generates two profunctors. One of the same direction $F_{*}:N\nrightarrow X$ and one of the opposite direction $F^{*}:X\nrightarrow N$. They are defined with the aid of hom-objects, where $F_{*}(x,i)=X(x,Fi)$ and $F^{*}(i,x)=X(Fi,x)$. In the case of Lawvere metric spaces, the profunctors of $F$ evaluated on objects $x$ and $i$ can be read as: ”The distance between $x$ and the image of $i$ under $F$”.